the sum of twice a number adn 7 less than the number is the same as the difference between -31 and the number. what is the number?
Answer by ewatrrr(24785)

 
Hi,
Let x represent the number
Question states***
2x+ (x-7) = (-31-x)
Solving for x
4x = -24
x = -6
CHECKING our Answer***
-12 -13 = -25 = -31-(-6) = -25